My question is: The dimensions of a triangle are multiplied by 1/4. The area of a smaller triangle can be found by multiplying the area of the original triangle by what number?
It seems too simple to say the answer is 1/4, but that's the only thing that makes sense to me. This is for 6th grade math. Thank you for any help you can provide!

Answer by rothauserc(4718)

let's pick a 3, 4, 5 right triangle, we know that
Area(A) of triangle = (1/2) * base * height
In the case of our triangle,
A = (1/2) * 3 * 4 = 6
now if we divide the dimensions by 4
A = (1/2) * (3/4) * (4/4) = (3/8)
so we want to solve,
6/x = 3/8
cross multiply fractions
3x = 48
x = 16
therefore,
we multiply the area of the larger triangle by (1/16)
***************************************************
lets check our rule with an isosceles right triangle which has sides 1, 1, sqrt(2)
A = (1/2) * 1 * 1 = (1/2)
also
A of smaller triangle = (1/2) * (1/4) * (1/4) = (1/32)
so we see that if we multiply (1/2) * (1/16) our rule works for a 1, 1, sqrt(2) isosceles right triangle also
*******************************************************